What is the pH of a solution that has a hydroxide ion concentration of 0.80×10−10 M?Express your answer to two significant figures.
Answer #1

We have

hydroxide ion concentration

[OH-] = 0.80 \times 10-10 M

Now we know the formulas

pOH = -log[OH-]

and pH + pOH = 14

So

pOH = -log[OH-] = -log(0.80 \times 10-10) = 10.09

Now

pH + pOH = 14

then

pH = 14 - pOH = 14 - 10.09 = 3.9

So our answer is

pH = 3.9